OpenStudy (anonymous):

I need help! practical math. Given a standard 52 card deck what is the probability of 3 eights will be dealt in a row? Extra credit if you use both methods. Also How about the probability of when each of the cards found is replaced back into the deck. I'll draw what I know so you can get what I mean

OpenStudy (anonymous):



OpenStudy (anonymous):

It depends on where in the deal you are. At the beginning, there are 52 cards from which 1 of 4 eights can be chosen. If it is an eight, then there remain 51 cards and 1 of 3 must be chosen. If that is an eight, there remain 50 cards, and 1 of 2 eights must be chosen. This is the most unlikely scenario, with a probability of (4/52)(3/51)(2/50) = 0.00018. If you replace the eights, you have (4/52)(4/52)(4/52)=0.00046.

OpenStudy (31356):

Correct!

OpenStudy (anonymous):

Thank you.

OpenStudy (ybarrap):

@darkside3704 - Did @douglaswinslowcooper 's great explanation help you?

OpenStudy (anonymous):

Yeah I wrote out the problem wrong it was suppose to be 52 51 50 lmfao. But yes, just the whole fractions that got me,

OpenStudy (anonymous):

How do you do you multiply fractions?

OpenStudy (ybarrap):

$$ \cfrac{2}{3}\times\cfrac{8}{5}=\cfrac{2\times 8}{3\times 5} $$

OpenStudy (anonymous):

(a/b)(c/d) = (ac)/(bd)